If y varies directly as x, and y = 16 when x = 20.
Find y when x = 5.

Answer by stanbon(75887)

If y varies directly as x, and y = 16 when x = 20.
Find y when x = 5.
---------
If y varies directl as x, y=kx; need to find "k".
16=k(20)
k=4/5
---------
EQUATION:
y=(4/5)x
If x=5 then y=(4/5)(5)=4
Cheers,
Stan H.